# Characterising and tackling thermally induced zero-drift in displacement measuring interferometry using temperature-controlled enclosure

### v1

@@ Abstract
Our research efforts in displacement measurement interferometry focused on long-term drifts initiated an extended experimental investigation in the interferometric assemblies of our design. We aimed to analyze, characterize and tackle the long-term measurement stability, expressed as the zero-drift, with special attention to the thermal effects. For the experimentation, we developed a thermostatic chamber equipped with an active temperature regulation, array of sensors and control electronics. With either the finely stabilized temperature or with the thermal cycling, we are able to carry out a range of investigations: verification of modified design or prototype interferometers, testing of production pieces, characterization of integrated assemblies and units in terms of the zero drift and the susceptibility to thermal effect - the temperature sensitivity, expressed as in nm.K.
@@ 1 Introduction
The extended time-frame examples typically occur in applications that require precise and stable positioning (large area scans in nanometrology, long expositions in e-beam lithography, reference object positioning in optical metrology), in applications where the interferometer(s) serve as a length reference (large volume metrology, calibrations with larger than usual number of measurement points or repetitions) and also in fundamental research (e.g. research on gravitational waves [5], realization of the unit kilogram using Kibble balance [6]).
The other part seeks to deliver the precision, inherent traceability and dynamic range of interferometric measurements outside the laboratories with a well-controlled and monitored environment, primarily temperature, air pressure and relative humidity. The applications range from 100 characterization of sensors elements [7] in production (as required by Industry 4.0), via fast turnover in-house calibration of working standards (e.g. gauge blocks, end bars) to in-situ calibration of instrumentation for a highly specific task (such as structural inspection in nuclear power plants [8]).
Our work aims to investigate, reduce, and characterize the residuals of zero-drifts, especially thermally induced, to make the unprecedented accuracy of the interferometric measurement applicable to these kinds of applications and measurement scenarios.
In this paper, we explore the length drifts beginning with the perspective of the entire measurement system assembly, quickly moving towards the system’s heart, the interferometer itself. During the journey, we identify the main culprit: the omnipresent effects of thermal changes on both the geometrical and material characteristics (Section 2). With this in mind, we design and realize a dedicated temperature-controlled enclosure as an experimental testbed to investigate the interferometric components and assemblies under thermal load (Section 3). We present a variety of (currently ongoing) investigations and an example of the improvement in tackling the zero drift in interferometric systems (Section 4). The paper is then summarised and concluded with future prospects in what we see as a broad potential for progress in interferometry-based dimensional metrology (Section 5).
@@ 2 Thermal effects influencing interferometric measurement
Temperature and its fluctuations profoundly influence dimensional measurement and impact the measurements’ accuracy and reliability. This influence primarily stems from the thermal expansion, which alters the geometry of both the measuring arrangement (not necessarily interferometric at this point) and the measurand. While the expansion of the measurand is typically accounted for (e.g. in gauge-block calibration[9]), accounting for the dimensional changes that influence the structural loop and, consequently, the measurement loop of the measurement systems can be challenging and is often neglected. The most straightforward way to limit the thermal effects at the system level is to keep the outer environment stable (a remarkable approach could be found in Helsinki [10]) and let the measurement equipment stabilize well before the actual measurement. The disadvantage is that any local temperature disturbances (e.g. heat dissipation from electronics or translation mechanisms) interfere with the temperature control and induce uneven temperature distribution along the measurement assembly. Another (complementary rather than alternative) approach to mitigate the thermal issues incorporates careful design of the measuring system focused on a well-stabilized metrology loop so that the thermal creeps have limited influence on the measurement arrangement. Typical examples could be using low-expansion materials for the metrology frame [11], using differential interferometers [12, 13] to shorten the loop and, generally, sticking to good practice for precision engineering [14].
@@
With interferometer(s) involved, the temperature has a massive influence on the refractive index of air. Unless the entire measurement system is put into vacuum [15], the measurement accuracy is hampered by K [16]. This influence is usually mitigated using a suitable empirical equation (i.e. indirect refractometry[17]). Nonetheless, the residual issues with a non-uniform temperature distribution of unknown magnitude could remain (for the comparison, the climate chambers used for such purposes usually declare the temperature homogeneity somewhere at C).
@@
A further element of interest is the susceptibility to the thermal effects of the interferometer itself, particularly the interferometer’s optics. Of the two mentioned influences – the thermally induced geometry changes and changes in the refractive index – the interferometer is significantly subject to both. With a temperature variation, the geometrical dimensions change, and so do the refractive index of the optical material the interferometer’s optics is manufactured of. The variation of the refractive index with temperature is referred to as the thermo-optic coefficient [18, 19]. Table 1 presents selected optical materials and corresponding values of the coefficient of the thermal expansion (CTE) and the thermo-optic coefficient (dn/dT [19]). The values indicate that these influences have a magnitude comparable to that of the thermal expansion and the fluctuations of the refractive index of air. To minimize the thermal drifts, it is crucial to design the optical arrangement of the interferometer in a way the parts of the interferometer’s arm within the interferometer’s optics are of equal optical length[20] and ideally share the path to a maximal extent. Despite these findings being decades old, it seems many state-of-the-art interferometric designs tend to neglect this aspect [21, 13]. Also, the thermal drifts were not thoroughly experimentally characterized in detail at the time of original investigation.
@@
CTE dn/dT Material (K) (K) Notes Schott N-BK7 7.1 1.29 Schott N-SF14 9.41 -1.46 Fused silica 0.55 12.9 Crystal quartz 7.5 -5.5 p-polarization 13.5 -6.5 s-polarization
@@
Finally, the technique of optical / opto-mechanical assembly can also render the interferometer sensitive to temperature changes. An interferometer assembled from individual components (rather than cemented or contacted together [13]) could be potentially susceptible to thermal drifts due to uneven influence of local temperature or pressure change or some misalignments that would, e.g. not exhibit on the fringe contrast directly. This susceptibility will differ depending on the particular arrangement and vary piece by piece due to the manufacturing process.
@@
In a real experiment or application, all the temperature-induced effect sum up and the resulting superpositions exhibits as the zero-drift. Due to non-ideal aspects of the reality, every compensation mechanism has its limits. It is inevitable that any interferometric measurement will suffer certain amount of residual thermal drift and exhibit some level of (dominantly thermally-induced) zero-drifts.
@@
These residual drifts become more pronounced in the scenarios mentioned above with the measurement over a longer period of time and in an environment with sub-optimal thermal conditions. The challenge we accepted was to specify and investigate a methodology allowing us to experimentally test the components of interferometric systems in an environment with stable and controlled temperature. This approach is intended to characterize the residual thermal drifts and quantify them as a measure called the temperature sensitivity, expressed as a displacement depending on the temperature in .
@@ 3 Temperature-controlled experimental enclosure
Typical instrumentation used to achieve temperature control in an air environment is some kind of climate chamber. Besides the costs, the major disadvantage is the homogeneity of the temperature inside the box (usually ) in conjunction with active air circulation, which is generally not favourable for interferometric measurements.
@@ 3.1 Enclosure design
To overcome these limitations so that we can characterize the thermal sensitivity we have designed a temperature-controlled experimental enclosure with the goal to achieve thermal stability better than mK. This (intermittent) goal was chosen so that the relative accuracy of displacement measurement, burdened with the thermal expansion and fluctuations of the air refractive index, would be at a similar level as the other significant source of drifts: the frequency source ( with a frequency-stabilized He-Ne laser). The entire ecosystem, sketched in Figure 1, consists of the actual enclosure, set of environmental sensors and control electronics.
The walls of the enclosure with a volume of litres and are lined with a polyethene insulation foam to provide the passive insulation. The enclosure’s floor holds a breadboard (for versatile assembly of tested devices), and its honeycomb construction ensures mechanical stability during thermal loads. An array of thermo-electric cells (six pieces) are mounted on the enclosure ceiling. One side is thermally connected to passive coolers outside the box, and the other to a homogenization plate inside. The enclosure is designed to be operated with the temperature set slightly below the room temperature so that the peltier cells cool down the inner space from the top, relying on thermal convection.
The chamber is equipped with a range of sensors (see also Figure 3). Several temperature sensors monitor the temperature conditions. The servo temperature sensor (T) serves as a process variable for the servo-loop temperature control. The reference temperature (T) sensor with calibrated absolute scale monitors the temperature at the same point as the servo-temperature sensor. The other pair monitors te external temperature (T) outside the enclosure, and the gradient temperature (T) is placed above the monitor and servo sensors (vertically spaced mm ) to detect the vertical temperature gradient. Besides the temperature, the air’s barometric pressure and relative humidity are measured.
The control electronics comprise three integral modules. The temperature controller TC interrogates the servo temperature (with mK resolution and excellent relative stability) and drives the current to the thermo-electric cells (up to A). The interferometric modul IFM interrogates the quadrature sine-cosine output either from the homodyne optical receivers or similar interface for the case of commercial instrumentation tested in the chambers. The environmental sensor control unit E.I.U (originally an indirect refractometer[22]) handles the rest of the sensors: three temperature sensors with the resolution of mK were calibrated with K (range from C to C, ), the pressure sensor (resolution Pa, range from kPa to kPa, kPa) and sensor for the relative humidity (). The E.I.U is powered from separate module PWR. Due to precision limits, a separate reference barometer (Druck PACE 1000) with Pa precision was operated along the enclosure.
@@ 3.2 Thermal conditions in the enclosure
To obtain insight into the thermal conditions in the enclosure, we analyzed two weeks long recording from the temperature sensors, displayed in Figure 2. The temperature fluctuated with mK as measured by both the servo temperature sensor and the reference temperature sensor. The good relative coincidence of the two sensors (mK), which is at the level of resolution of the less precise sensor, proves there are no long-term drifts in the temperature measurement. The recording of the external temperature (outside the box) indicates that the external temperature and inner temperatures are coupled (and further tuning of the servo-control is necessary). Finally, the measurement revealed a negative vertical temperature gradient inside the enclosure of approximately K.m, which is quite significant and requires further investigation.

### v2

@@ Abstract
Our research efforts in displacement measurement interferometry focused on long-term drifts initiated an extended experimental investigation in the interferometric assemblies of our design. We aimed to analyze, characterize and tackle the long-term measurement stability, expressed as the zero-drift, with particular attention to the thermal effects. For the experimentation, we developed a thermostatic chamber equipped with active temperature regulation, an array of sensors and control electronics. With either the finely stabilized temperature or with the thermal cycling, we can carry out a range of investigations: verification of modified design or prototype interferometers, testing of production pieces, characterization of integrated assemblies and units in terms of the zero drift and the susceptibility to thermal effects – the temperature sensitivity (, expressed in nm.K-1). With these experimental studies, we demonstrate the potential of the zero drift studies to contribute to the development and broader expansion of interferometric instrumentation.
@@
Published article: Simon Rerucha et al 2024, Meas. Sci. Technol. 35 115011
https://doi.org/10.1088/1361-6501/ad646b
@@ 1 Introduction
The extended time-frame examples typically occur in applications that require precise and stable positioning (large area scans in nanometrology, long expositions in e-beam lithography, reference object positioning in optical metrology), in applications where the interferometer(s) serve as a length reference (large volume metrology, calibrations with larger than the usual number of measurement points or repetitions) and also in fundamental research (e.g. research on gravitational waves [5], the realization of the unit kilogram using Kibble balance [6]).
The other part seeks to deliver the precision, inherent traceability, and dynamic range of interferometric measurements operated in a well-controlled and monitored environment (primarily the temperature, air pressure, and relative humidity) outside the laboratories. The applications range from 100 characterization of sensors elements [7] in production (as required by Industry 4.0) via fast turnover in-house calibration of working standards (e.g. gauge blocks, end bars) to in-situ calibration of instrumentation for highly specific tasks (such as structural inspection in nuclear power plants [8]).
Our work aims to investigate, characterize and possibly reduce zero-drifts, especially tge thermally induced ones, to make the unprecedented accuracy of the interferometric measurement applicable to these kinds of applications and measurement scenarios. Regarding the application parameters, our work generally targets the range between the micro- and nanoscale (broadly covered by the nanometrology community today) and the long-range measurements. We thus aim at the metre range, i.e. lengths between m and m, with a particular focus on the lower part of the range that matches e.g. with the coordinate positioning. With precision and accuracy, we are targeting the contribution to the measurement uncertainty of to reach a comparable level to the other fundamental limited, the influence of the refractive index of air. As the temperature-induced zero-drifts of an interferometer scale primarily with the arrangement and dimensions of the optical part and the magnitude of the temperature fluctuations rather than the measured distance, we practically aim at a nm.K target.
In this paper, we explore the length drifts beginning with the perspective of the entire measurement system assembly, quickly moving towards the system’s heart, the interferometer itself. During the journey, we identify the main culprit: the omnipresent effects of thermal changes on both the geometrical and material characteristics (Section 2). With this in mind, we design and realize a dedicated temperature-controlled enclosure as an experimental testbed to investigate the interferometric components’ and assemblies’ temperature sensitivity under thermal load (Section 3). Then we present four experimental studies, demonstrating how the zero drift studies can contribute to the characterization, verification and tackling the zero drift in interferometric systems (Section 4). The paper is then summarised and concluded with future prospects in what we see as a broad potential for progress in interferometry-based dimensional metrology (Section 5).
@@ 2 Thermal effects influencing interferometric measurement
As it is well known, temperature and its fluctuations profoundly influence dimensional measurement and impact the measurements’ accuracy and reliability. With laser interferometry, the fluctuating temperature influences all components of any measurement assembly – laser source (and delivery), the interferometer and optics, the (opto-) mechanics, the environment (air) and detection electronics – in a way that might exhibit as thermally induced zero drifts [9].
This influence primarily stems from the thermal expansion, which alters the geometry of both the measuring arrangement (not necessarily interferometric at this point) and the measurand. While the expansion of the measurand is typically accounted for (e.g. in gauge-block calibration [10]), accounting for the dimensional changes that influence the structural loop and, consequently, the measurement loop of the measurement systems can be challenging and is often neglected.
@@
The most straightforward way to limit the thermal effects at the system level is to keep the outer environment stable (a remarkable approach could be found in Helsinki [11]) and let the measurement equipment stabilize well before the actual measurement. The disadvantage is that any local temperature disturbances (e.g. heat dissipation from electronics or translation mechanisms) interfere with the temperature control and induce uneven temperature distribution along the measurement assembly. Another (complementary rather than alternative) approach to mitigating thermal issues incorporates a careful design of the measuring system focused on a well-stabilized metrology loop so that thermal creeps have a limited influence on the measurement arrangement. Typical examples could be using low-expansion materials for the metrology frame [12], using differential interferometers [13, 14] to shorten the loop and, generally, sticking to good practice for precision engineering [15].
@@
With interferometer(s) involved, the temperature also has a massive influence on the refractive index of air, as summarized in Table LABEL:t:ria. Unless the entire measurement system is put into vacuum [16], the measurement accuracy is hampered by K-1 [17]. This influence is usually mitigated using a suitable empirical equation (i.e. indirect refractometry [18]). Nonetheless, the residual issues with a non-uniform temperature distribution of unknown magnitude could remain (for the comparison, the climate chambers used for such purposes usually declare the temperature homogeneity somewhere at C). For practical applications, it is, however, essential to consider the degree of precision needed for measurement of the other parameters of the air, namely the atmospheric pressure, to reduce both absolute errors and potential drifts (see the fourth column in Table LABEL:t:ria that states the relative error contribution stemming from a 1% measurement error).
@@
A further element of interest is the susceptibility to the thermal effects of the interferometer itself, particularly the interferometer’s optics [9]. With fluctuating temperatures, the geometrical dimensions change, and so does the refractive index of the optical material the interferometer’s optics is manufactured of. The variation of the refractive index with temperature is referred to as the thermo-optic coefficient [19, 20]. Table LABEL:t:properties presents selected optical materials and corresponding values of the coefficient of the thermal expansion (CTE) and the thermo-optic coefficient (dn/dT [20]). The values indicate that these influences might have a magnitude comparable to that of the thermal expansion and the fluctuations of the refractive index of air; nonetheless, the magnitudes are not comparable directly, as the refraction influence scales with the dead-path in the air, the thermal drift of the optical system scales with its dimensions.
To minimize the thermal drifts, it is crucial to design the optical arrangement of the interferometer in a way the parts of the interferometer’s arm within the interferometer’s optics are of equal optical length[21, 9] and ideally share the path to a maximal extent. Despite these decades-old findings, it seems many state-of-the-art interferometric designs tend to neglect this aspect; examples to be found, e.g. in [22, 14]. Also, to our best knowledge, the thermal drifts were not thoroughly experimentally characterized in detail at the time of the original investigation.
@@
CTE dn/dT Material (K) (K) Notes Schott N-BK7 7,1 1,29 Schott N-SF14 9,41 -1,46 Fused silica 0,55 12,9 Crystal quartz 7,5 -5,5 perpendicular to opt. axis 13,5 -6,5 parallel to opt. axis Refractive index of air – 0,95
Value at standard Refractive index R. I. change per Influence parameter conditions sensitivity coefficient value change Temperature C K-1 Pressure Pa Pa-1 Relative humidity % (% RH))-1 CO2 content ppm (ppm))-1
@@
Finally, the optical / opto-mechanical assembly technique can also render the interferometer sensitive to temperature changes. An interferometer assembled from individual components (rather than cemented or contacted together [14]) could be potentially susceptible to thermal drifts due to uneven influence of local temperature or pressure change or some misalignments that would, e.g. not exhibit on the fringe contrast directly. This susceptibility will differ depending on the particular arrangement and vary piece by piece due to the manufacturing process.
@@
In a real-world experiment or application, all the temperature-induced effects sum up, and the resulting superposition contributes to the zero drift. Due to non-ideal aspects of reality, every compensation mechanism has its limits. Any interferometric measurement will inevitably suffer a certain amount of residual thermal drift and exhibit some level of (dominantly thermally-induced) zero-drifts. These residual drifts become more pronounced in the scenarios mentioned above with the measurement over a more extended period of time and in an environment with sub-optimal thermal conditions.
@@
The challenge we accepted was to specify and investigate a methodology allowing us to experimentally test the components of interferometric systems in an environment with stable and controlled temperatures. This approach is intended to characterize both the systematic and residual thermal drifts and quantify them as a measure called the temperature sensitivity, expressed as a displacement depending on the temperature in .K-1.
The quantitative knowledge of this value is directly applicable as an input to the uncertainty analysis when the interferometric system is integrated into a measurement instrument. It is also helpful for the iterative development or improvement of either a single component (subsystem) or gradually extended assemblies (e.g. from standalone interferometer’s optics to integrated interferometric assembly including fixtures), where it is possible to evaluate the gradual increase (desirably minimal) of the temperature sensitivity. In this case, even a quantitative estimation of the thermal sensitivity could aid the process, allowing for side-by-side comparison. Last but not least, the established (and experimentally verified) knowledge can serve as a reference figure for the verification and testing in the manufacturing and production of the interferometers.
@@ 3 Temperature-controlled experimental enclosure
Typical instrumentation used to achieve temperature control in an air environment is some kind of climate chamber. Besides the costs, the major disadvantage is the inhomogeneity of the temperature inside the box in conjunction with forced air circulation, which is generally not favourable for interferometric measurements.
To overcome these limitations and allow for the characterization of the thermal sensitivity of interferometric systems, we have designed a temperature-controlled experimental enclosure to achieve thermal stability better than mK. This (intermittent) goal was chosen so that the relative accuracy of displacement measurement, burdened with the thermal expansion and fluctuations of the air refractive index, would be at a similar level as the other significant source of drifts: the frequency source ( with a frequency-stabilized He-Ne laser). In the following sections, we briefly present the design and construction of the enclosure, the achievable thermal condition within the enclosure, and the methodology we used for the experimentation.
@@ 3.1 Enclosure design
The entire ecosystem, sketched in Figure 1, consists of the actual enclosure, set of environmental sensors and control electronics. The enclosure with a volume of litres is designed to accomodate a 300mm x 600mm optical breadboard. The walls (mm Birch plywood) are lined with a polyethene insulation foam (mm plates) to provide the passive insulation. The enclosure’s floor holds the breadboard (for versatile assembly of tested devices), and its honeycomb construction ensures mechanical stability during thermal loads. An array of thermo-electric cells (six pieces) are mounted on the enclosure ceiling. One side is thermally connected to passive coolers outside the box, and the other to a homogenization plate inside (mm aluminium). The enclosure is designed to be operated with the temperature set slightly below the room temperature as we aim to limit the temperature difference between the chamber volume and the outside temperature and reduce the vertical temperature gradient in the chamber. This way, the warmer air rises upwards (natural convection), where the Peltier cells dissipate the heat out of the chamber, and the cooler air sinks down.
The chamber is equipped with a range of sensors (see also Figure 3). Several temperature sensors monitor the temperature conditions. The servo temperature sensor () serves as a process variable for the servo-loop temperature control. The reference temperature () sensor with a calibrated absolute scale monitors the temperature at the same point as the servo-temperature sensor. The other pair monitors the external temperature () outside the enclosure, and the gradient temperature () is placed above the monitor and servo sensors (with defined vertical spacing ) to detect the vertical temperature gradient. Besides the temperature, the air’s barometric pressure and relative humidity are measured.
The control electronics comprise three integral modules. The temperature controller TC interrogates the servo temperature (with mK resolution) and drives the current to the thermo-electric cells (up to A). The interferometric modul IFM interrogates the quadrature sine-cosine output either from the homodyne optical receivers or similar interface for the case of commercial instrumentation tested in the chambers. The environmental sensor interrogator unit E.I.U (originally an indirect refractometer [23]) handles the rest of the sensors: three temperature sensors (Analog Devices AD22100) with the resolution of mK were calibrated with K (range from C to C, ), the pressure sensor (NXP MPXH6101A; resolution Pa, range from kPa to kPa, kPa) and sensor for the relative humidity (Humirel HTM2500; ). A separate module, PWR, powers the E.I.U. Due to precision limits, a separate reference barometer (Druck PACE 1000) with Pa precision was operated along the enclosure.
@@ 3.2 Thermal conditions in the enclosure
To obtain insight into the thermal conditions in the enclosure, we analyzed day-long recording from the temperature sensors, displayed in Figure 2. The temperature fluctuated with mK as measured by both the servo temperature sensor and the reference temperature sensor. The good relative coincidence of the two sensors (mK), which is at the level of resolution of the less precise sensor, proves there are no long-term drifts in the temperature measurement. The recording of the external temperature (outside the box) indicates that the external and inner temperatures are coupled (and further tuning of the servo-control is necessary). Finally, the measurement revealed a negative vertical temperature gradient inside the enclosure of approximately K.m-1, which is quite significant and requires further investigation, as well as the homogeneity within the enclosure in general. Preliminary experiments carried out with a smaller version of the enclosure ( litre volume with mm insulation) indicate achievable temperature stability below mK over a week and the horizontal gradient below K.
3.3 Typical experimental test-bed and methodology
An outcome of preliminary testing (not covered in this paper) is a somehow standardized experimental arrangement and protocol for the experimentation. In a typical arrangement, depicted in Figure 3, the interferometer was adjusted against a single, fixed mirror, keeping the separation between the interferometers’ front side and the mirror minimal.
The servo-controlled temperature in the enclosure allowed for two types of investigation: (i) the temperature is kept stabilized (actively or only passively) for a prolonged period (days to weeks) so that the limits of the thermo-mechanical stability of individual components or more complex assemblies can be investigated, or (ii) the temperature was modulated (and thus the thermal load applied) in predefined cycles with linear ramp-up.
With the bulk interferometers, the feasible temperature change rate was experimentally identified to be between Khr-1 and Khr-1. This value represents a trade-off between the quite a small power rating of the cooling relative to the enclosure volume plus the need to let all components change the temperature as homogeneously as possible and the need to conduct the experiment in a timely manner. The preliminary experimentation revealed that higher heating/cooling rate led to more destabilized thermal conditions in the enclosure and, in turn, distorted measurement data (e.g. significant presence of overshoots in the displacement indications, apparent phase delay between the temperature modulation and displacement response, noticeable hysteresis in the correlation plot). An open question for future investigation is using harmonic temperature modulation; this approach has the potential to reduce the experimental time, but the risky aspect is that the thermal response of the investigated system might get attenuated with higher modulation frequencies and that the attenuation might not be as apparent as it is now with the trapezoidal modulation profile.
The temperature sensors were placed on a pillar post close to the center of the enclosure, with the servo and reference sensor being but to the same height as the interferometers’ optics and the gradient sensor above them (mm).
The laser beam was delivered using a polarization-maintaining fibre either from a frequency-stabilized He-Ne laser or from an iodine-stabilized semiconductor laser system [24]. Any residual fluctuation in the incoming beam polarization was filtered out with a calcite polarizer (e.g. Glan-Thompson type). When suitable, the interferograms were processed with a four-detector type homodyne receiver (shown in Figure 5 d) and described in [25]). The symmetric nature of this receiver architecture provides good resilience against amplitude fluctuations of the interferogram signal. The interferometer-mirror-receiver triplet(s) were adjusted with the help of a digital oscilloscope to obtain a symmetric and well-centred Lissajeous diagram at the receiver’s output. The analogue output of the receiver is digitized with the IFM module. Any residual offset in the analogue-to-digital conversion is calibrated before the measurement, and the drifts of the offset are negligible.
The data from the homodyne receivers, control modules and sensors was recorded using a unified communication framework (CANopen protocol with custom extension) at a native sampling rate of individual modules (between Hz and Hz), subsampled to Hz and evaluated. Before the actual test run, after the interferometers were adjusted, we manually induced vibrations to the system and simultaneously recorded the response to (i) double-check the counting direction and (ii) obtain data for elliptic correction [26], which was later applied to the recorded data to remove any remaining offsets or asymmetry in the quadrature output; example recording is shown in Figure 4. The follow-up data analysis incorporated the removal of a residual linear drift, applying the linearity correction and the refractive index of air compensation, evaluating the correlation between the induced temperature modulation and drifts in the interferometers’ displacement reading, and final interpretation.
The total relative uncertainty associated with the refractive index measurement was (), so with the distance between the front face of the interferometer and the mirror mm, the absolute uncertainty was nm. A significantly larger contribution could be expected from the thermal expansion of the experimental optomechanics, as discussed later.